# Hardness of Range Avoidance and

### v1

@@ 2.4 Proof Complexity
Recall that $\mathrm{TAUT}$}, the set of DNF tautologies, is the canonical $\mathbf{coNP}$}-complete problem. A propositional proof system is simply a nondeterministic algorithm for $\mathrm{TAUT}$}. More formally:
@@
  • (Completeness) For every $\varphi\in{\mathrm{TAUT}}$, there exists a string $\pi\in\{0,1\}^{*}$ such that $\mathcal{P}(\varphi,\pi)$ accepts.
  • (Soundness) For every $\varphi,\pi\in\{0,1\}^{*}$, if $\mathcal{P}(\varphi,\pi)$ accepts, then $\varphi\in{\mathrm{TAUT}}$.
@@
The range avoidance problem with suitable stretch is in $\mathbf{FNP}$} if and only if there exists a propositional proof system that breaks every proof complexity generator.
@@ 3 Hardness of Range Avoidance
Assume that for some $m>n$, there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ and $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ is an $(N-1,2^{-m-1})$-strong seeded extractor. ($N,d\leq\mathrm{poly}(m)$.) Then Avoid for polynomial-size circuits of stretch $n\to m$ is not in $\mathbf{SearchNP}$}.
@@ 4.1 Separating $\mathsf{PV}_{1}$ from $\mathsf{APC}_{1}$
As discussed in Section 2.5, to separate $\mathsf{PV}_{1}$ from $\mathsf{APC}_{1}$, it suffices to show that there is no polynomial-time Student that wins the Student-Teacher game for Avoid in $O(1)$ rounds. In fact, we will show something stronger: Let $k=k(n)$ be a parameter, assuming the existence of demi-bits generators secure against ${\mathbf{AM}}/_{O(\log k)}$, there is no polynomial-time Student that wins the Student-Teacher game for Avoid in $k(n)$ rounds.
@@
Let $m,n,k$ be parameters such that $m>n$. Assume there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ secure against ${\mathbf{AM}}/_{O(\log k)}$. Let $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be an $(N-1,2^{-10km})$-strong extractor. Then for every deterministic polynomial-time Student $A$, there is a string $r\in\{0,1\}^{d}$ and a Teacher such that $A$ fails to solve Avoid on $C_{r}$ in $k$ rounds, where $C_{r}:\{0,1\}^{n}\to\{0,1\}^{m}$ is the circuit
@@
Consider the following ${\mathbf{AM}}/_{O(\log k)}$ protocol that attempts to break the demi-bits generator $G$. This protocol has the index $j$ hardwired as advice but is otherwise uniform.
@@ 2 $\mathsf{Prover}$ sends $s_{1},\dots,s_{j-1}$;
Algorithm 4.1 The ${\mathbf{AM}}/_{O(\log k)}$ protocol $\mathcal{P}$ breaking demi-bits generator $G$
@@
Since $G$ is a demi-bits generator secure against ${\mathbf{AM}}/_{O(\log k)}$ adversaries, $\mathcal{P}$ does not have the soundness for all sufficiently large $n$. In other words, there is a $\mathsf{Prover}^{*}$ that makes the $\mathsf{Verifier}$ accepts some $y\in\mathrm{Range}(G)$ with probability >1/3>1/3. Fix such a string $y$, let $s_{j}$ be any $n$-bit string such that $G(s_{j})=y$, and let $s_{1},\dots,s_{j-1}\in\{0,1\}^{n}$ be the message sent in Algorithm 4.1 (of Algorithm 4.1) by $\mathsf{Prover}^{*}$ on the input $y$.
@@ 4.2 From Demi-Bits to Pseudo-Surjectivity
Let $G:\{0,1\}^{n}\to\{0,1\}^{N}$ be a demi-bits generator secure against ${\mathbf{NP}}/_{\mathrm{poly}}$, $k\in\mathbb{N}$, and $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be an $(N-1,\varepsilon)$-strong linear extractor for $\varepsilon:=2^{-10km}$. ($k,d,N\leq\mathrm{poly}(m)$.) Then for every non-uniform propositional proof system $\mathcal{P}$, there is a string $r\in\{0,1\}^{d}$ such that the circuit $C_{r}:\{0,1\}^{n}\to\{0,1\}^{m}$ defined as
@@
Suppose for any parameter $N\leq\mathrm{poly}(n)$, there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ secure against ${\mathbf{NP}}/_{\mathrm{poly}}$. Then for every non-uniform propositional proof system $\mathcal{P}$ and any parameters $k\leq\mathrm{poly}(n)$ and $n<m<\mathrm{poly}(n)$, there is a circuit $C:\{0,1\}^{n}\to\{0,1\}^{m}$ of size $\mathrm{poly}(n)$ such that $C$ is a $k$-round pseudo-surjective proof complexity generator secure against $\mathcal{P}$.
@@ 5 Hardness of Range Avoidance from Predictable Arguments
Ilango, Li, and Williams [ILW23] proved ${\mathrm{\textnormal{{Avoid}}}}\not\in{\mathbf{FP}}$ assuming the existence of JLS-secure $i\mathcal{O}$ and ${\mathbf{NP}}\neq{\mathbf{coNP}}$. Given our main result that the existence of demi-bits generators implies the hardness of Avoid, it is natural to ask whether the $i\mathcal{O}$ assumption used in [ILW23] can be weakened to a “Minicrypt” assumption. In particular, we conjecture:
@@
${\mathrm{\textnormal{{Avoid}}}}\not\in{\mathbf{FP}}$ follows from the existence of one-way functions and ${\mathbf{NP}}\neq{\mathbf{coNP}}$.
Unfortunately, we are not able to prove 5.1. Instead, in this section, we present an alternative interpretation of [ILW23]. Although the results and proofs are not new, we hope that our new perspective helps make progress towards proving 5.1, or in general, basing ${\mathrm{\textnormal{{Avoid}}}}\not\in{\mathbf{FP}}$ from minimal assumptions.
@@
Instead of $i\mathcal{O}$, what [ILW23] actually needs is witness encryption [GGSW13]. Let $L\in{\mathbf{NP}}$ (usually we take $L$ to be an $\mathbf{NP}$}-complete language such as $\mathrm{SAT}$}). A witness encryption scheme for $L$ is a pair of algorithms $(\mathsf{Enc},\mathsf{Dec})$ with the following interface: [^9] Witness encryption schemes as defined in [GGSW13] can only encrypt a one-bit message $m\in\{0,1\}$. To obtain a witness encryption scheme that can encrypt arbitrarily long message, one can simply encrypt each message bit independently. The security of this new scheme follows easily from a hybrid argument.
@@
Now, suppose that there is a deterministic algorithm $\mathcal{A}$ solving Avoid, we show that $L\in{\mathbf{coNP}}$. The idea is to use $\mathcal{A}$ as the prover in the above argument system. In particular, let $C_{x,ct}:\{0,1\}^{m}\to\{0,1\}^{\ell}$ denote the circuit that given $w\in\{0,1\}^{m}$ as input, outputs $\mathsf{Dec}(x,w,ct,1^{\lambda})$. Upon receiving $ct$, $\mathsf{Prover}$ always sends $\mathcal{A}(C_{x,ct})$ to $\mathsf{Verifier}$.
@@
In summary, witness encryption implies that $\mathbf{NP}$} has a special type of predictable argument system. Moreover, if ${\mathrm{\textnormal{{Avoid}}}}\in{\mathbf{FP}}$, then plugging the Avoid algorithm as the prover results in the following intriguing situation: if $x\in L$, then the $\mathsf{Verifier}$ never accepts, while if $x\not\in L$, then the $\mathsf{Verifier}$ accepts with non-zero probability! This allows us to put every language with such argument systems in $\mathbf{coNP}$}.
An interesting question is to identify the weakest possible argument system for $\mathbf{NP}$} such that the above situation happens. Can we build such argument systems using only one-way functions? Such an argument system would make progress towards proving 5.1.
@@ Appendix A Ilango’s Proof
We present an alternative proof of our main results that demi-bits generators imply ${\mathrm{\textnormal{{Avoid}}}}\not\in{\mathbf{SearchNP}}$. This proof is due to Rahul Ilango [Ila25] in an independent and concurrent work.
@@
The proof relies on the following fact underlying Lautemann’s proof [Lau83] that $\mathbf{BPP}$} is in the polynomial hierarchy.
@@ B.1 Demi-Bits Generators with Polynomial Stretch
For a Boolean function $f:\{0,1\}^{n}\to\{0,1\}$ and a parameter $s(n)\leq\mathrm{poly}(n)$, let $\mathsf{lb}(f,s)$ denote the sentence stating that “$f$ requires circuit complexity at least s(n)s(n)”. This sentence can be written as a CNF of size $2^{O(n)}$ and, if true, admits a trivial proof of length $2^{\tilde{O}(s(n))}$ in most reasonable proof systems. Rudich’s conjecture asserts that there is no (non-uniform) propositional proof system that has length-$2^{O(n)}$ proofs of $\mathsf{lb}(f,s)$ for a large fraction of Boolean functions $f$. This can be equivalently formulated as the non-existence of ${\mathbf{NP}}/_{\mathrm{poly}}$-natural proofs” against polynomial-size circuits (see [Rud97] for more details). Rudich’s conjecture was further investigated in [PS19, ST21].
@@ B.3 Constant-Degree Demi-Bits Generators from Goldreich’s Generator
is an influential candidate pseudorandom generator with large stretch that is computable with constant locality (i.e., in ${\mathbf{NC}}^{0}$[Gol11]. To define the generator $G:\{0,1\}^{n}\to\{0,1\}^{m}$, fix a $d$-uniform hypergraph with $n$ vertices and $m$ (ordered) hyperedges each of size $d$, and a predicate $P:\{0,1\}^{d}\to\{0,1\}$. For each $i\in[m]$, the $i$-th output bit is obtained by applying $P$ to the input bits on the $i$-th hyperedge. That is, suppose the $i$-th hyperedge contains vertices $v_{i,1},v_{i,2},\dots,v_{i,d}$, then on input $x\in\{0,1\}^{n}$, the $i$-th output bit is

### v2

@@ 2.4 Proof Complexity
Recall that $\mathsf{TAUT}$}, the set of DNF tautologies, is the canonical $\mathsf{coNP}$}-complete problem. A propositional proof system is simply a nondeterministic algorithm for $\mathsf{TAUT}$}. More formally:
@@
  • (Completeness) For every $\varphi\in{\mathsf{TAUT}}$, there exists a string $\pi\in\{0,1\}^{*}$ such that $\mathcal{P}(\varphi,\pi)$ accepts.
  • (Soundness) For every $\varphi,\pi\in\{0,1\}^{*}$, if $\mathcal{P}(\varphi,\pi)$ accepts, then $\varphi\in{\mathsf{TAUT}}$.
@@
The range avoidance problem with suitable stretch is in $\mathsf{FNP}$} if and only if there exists a propositional proof system that breaks every proof complexity generator.
@@ 3 Hardness of Range Avoidance
Assume that for some $m>n$, there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ and $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ is an $(N-1,2^{-m-1})$-strong seeded extractor. ($N,d\leq\mathrm{poly}(m)$.) Then Avoid for polynomial-size circuits of stretch $n\to m$ is not in $\mathsf{SearchNP}$}.
@@ 4.1 Separating $\mathsf{PV}_{1}$ from $\mathsf{APC}_{1}$
As discussed in Section 2.5, to separate $\mathsf{PV}_{1}$ from $\mathsf{APC}_{1}$, it suffices to show that there is no polynomial-time Student that wins the Student-Teacher game for Avoid in $O(1)$ rounds. In fact, we will show something stronger: Let $k=k(n)$ be a parameter, assuming the existence of demi-bits generators secure against ${\mathsf{AM}}/_{O(\log k)}$, there is no polynomial-time Student that wins the Student-Teacher game for Avoid in $k(n)$ rounds.
@@
Let $m,n,k$ be parameters such that $m>n$. Assume there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ secure against ${\mathsf{AM}}/_{O(\log k)}$. Let $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be an $(N-1,2^{-10km})$-strong extractor. Then for every deterministic polynomial-time Student $A$, there is a string $r\in\{0,1\}^{d}$ and a Teacher such that $A$ fails to solve Avoid on $C_{r}$ in $k$ rounds, where $C_{r}:\{0,1\}^{n}\to\{0,1\}^{m}$ is the circuit
@@
Consider the following ${\mathsf{AM}}/_{O(\log k)}$ protocol that attempts to break the demi-bits generator $G$. This protocol has the index $j$ hardwired as advice but is otherwise uniform.
@@ 2 $\mathsf{Prover}$ sends $s_{1},\dots,s_{j-1}$;
Algorithm 4.1 The ${\mathsf{AM}}/_{O(\log k)}$ protocol $\mathcal{P}$ breaking demi-bits generator $G$
@@
Since $G$ is a demi-bits generator secure against ${\mathsf{AM}}/_{O(\log k)}$ adversaries, $\mathcal{P}$ does not have the soundness for all sufficiently large $n$. In other words, there is a $\mathsf{Prover}^{*}$ that makes the $\mathsf{Verifier}$ accepts some $y\in\mathrm{Range}(G)$ with probability >1/3>1/3. Fix such a string $y$, let $s_{j}$ be any $n$-bit string such that $G(s_{j})=y$, and let $s_{1},\dots,s_{j-1}\in\{0,1\}^{n}$ be the message sent in Algorithm 4.1 (of Algorithm 4.1) by $\mathsf{Prover}^{*}$ on the input $y$.
@@ 4.2 From Demi-Bits to Pseudo-Surjectivity
Let $G:\{0,1\}^{n}\to\{0,1\}^{N}$ be a demi-bits generator secure against ${\mathsf{NP}}/_{\mathrm{poly}}$, $k\in\mathbb{N}$, and $\mathsf{Ext}:\{0,1\}^{N}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be an $(N-1,\varepsilon)$-strong linear extractor for $\varepsilon:=2^{-10km}$. ($k,d,N\leq\mathrm{poly}(m)$.) Then for every non-uniform propositional proof system $\mathcal{P}$, there is a string $r\in\{0,1\}^{d}$ such that the circuit $C_{r}:\{0,1\}^{n}\to\{0,1\}^{m}$ defined as
@@
Suppose for any parameter $N\leq\mathrm{poly}(n)$, there exists a demi-bits generator $G:\{0,1\}^{n}\to\{0,1\}^{N}$ secure against ${\mathsf{NP}}/_{\mathrm{poly}}$. Then for every non-uniform propositional proof system $\mathcal{P}$ and any parameters $k\leq\mathrm{poly}(n)$ and $n<m<\mathrm{poly}(n)$, there is a circuit $C:\{0,1\}^{n}\to\{0,1\}^{m}$ of size $\mathrm{poly}(n)$ such that $C$ is a $k$-round pseudo-surjective proof complexity generator secure against $\mathcal{P}$.
@@ 5 Hardness of Range Avoidance from Predictable Arguments
Ilango, Li, and Williams [ILW23] proved ${\mathsf{\textnormal{{Avoid}}}}\not\in{\mathsf{FP}}$ assuming the existence of JLS-secure $i\mathcal{O}$ and ${\mathsf{NP}}\neq{\mathsf{coNP}}$. Given our main result that the existence of demi-bits generators implies the hardness of Avoid, it is natural to ask whether the $i\mathcal{O}$ assumption used in [ILW23] can be weakened to a “Minicrypt” assumption. In particular, we conjecture:
@@
${\mathsf{\textnormal{{Avoid}}}}\not\in{\mathsf{FP}}$ follows from the existence of one-way functions and ${\mathsf{NP}}\neq{\mathsf{coNP}}$.
Unfortunately, we are not able to prove 5.1. Instead, in this section, we present an alternative interpretation of [ILW23]. Although the results and proofs are not new, we hope that our new perspective helps make progress towards proving 5.1, or in general, basing ${\mathsf{\textnormal{{Avoid}}}}\not\in{\mathsf{FP}}$ from minimal assumptions.
@@
Instead of $i\mathcal{O}$, what [ILW23] actually needs is witness encryption [GGSW13]. Let $L\in{\mathsf{NP}}$ (usually we take $L$ to be an $\mathsf{NP}$}-complete language such as $\mathsf{SAT}$}). A witness encryption scheme for $L$ is a pair of algorithms $(\mathsf{Enc},\mathsf{Dec})$ with the following interface: [^10] Witness encryption schemes as defined in [GGSW13] can only encrypt a one-bit message $m\in\{0,1\}$. To obtain a witness encryption scheme that can encrypt arbitrarily long message, one can simply encrypt each message bit independently. The security of this new scheme follows easily from a hybrid argument.
@@
Now, suppose that there is a deterministic algorithm $\mathcal{A}$ solving Avoid, we show that $L\in{\mathsf{coNP}}$. The idea is to use $\mathcal{A}$ as the prover in the above argument system. In particular, let $C_{x,ct}:\{0,1\}^{m}\to\{0,1\}^{\ell}$ denote the circuit that given $w\in\{0,1\}^{m}$ as input, outputs $\mathsf{Dec}(x,w,ct,1^{\lambda})$. Upon receiving $ct$, $\mathsf{Prover}$ always sends $\mathcal{A}(C_{x,ct})$ to $\mathsf{Verifier}$.
@@
In summary, witness encryption implies that $\mathsf{NP}$} has a special type of predictable argument system. Moreover, if ${\mathsf{\textnormal{{Avoid}}}}\in{\mathsf{FP}}$, then plugging the Avoid algorithm as the prover results in the following intriguing situation: if $x\in L$, then the $\mathsf{Verifier}$ never accepts, while if $x\not\in L$, then the $\mathsf{Verifier}$ accepts with non-zero probability! This allows us to put every language with such argument systems in $\mathsf{coNP}$}.
An interesting question is to identify the weakest possible argument system for $\mathsf{NP}$} such that the above situation happens. Can we build such argument systems using only one-way functions? Such an argument system would make progress towards proving 5.1.
@@ Appendix A Ilango’s Proof
We present an alternative proof of our main results that demi-bits generators imply ${\mathsf{\textnormal{{Avoid}}}}\not\in{\mathsf{SearchNP}}$. This proof is due to Rahul Ilango [Ila25] in an independent and concurrent work.
@@
The proof relies on the following fact underlying Lautemann’s proof [Lau83] that $\mathsf{BPP}$} is in the polynomial hierarchy.
@@ B.1 Demi-Bits Generators with Polynomial Stretch
For a Boolean function $f:\{0,1\}^{n}\to\{0,1\}$ and a parameter $s(n)\leq\mathrm{poly}(n)$, let $\mathsf{lb}(f,s)$ denote the sentence stating that “$f$ requires circuit complexity at least s(n)s(n)”. This sentence can be written as a CNF of size $2^{O(n)}$ and, if true, admits a trivial proof of length $2^{\tilde{O}(s(n))}$ in most reasonable proof systems. Rudich’s conjecture asserts that there is no (non-uniform) propositional proof system that has length-$2^{O(n)}$ proofs of $\mathsf{lb}(f,s)$ for a large fraction of Boolean functions $f$. This can be equivalently formulated as the non-existence of ${\mathsf{NP}}/_{\mathrm{poly}}$-natural proofs” against polynomial-size circuits (see [Rud97] for more details). Rudich’s conjecture was further investigated in [PS19, ST21].
@@ B.3 Constant-Degree Demi-Bits Generators from Goldreich’s Generator
is an influential candidate pseudorandom generator with large stretch that is computable with constant locality (i.e., in ${\mathsf{NC}}^{0}$[Gol11]. To define the generator $G:\{0,1\}^{n}\to\{0,1\}^{m}$, fix a $d$-uniform hypergraph with $n$ vertices and $m$ (ordered) hyperedges each of size $d$, and a predicate $P:\{0,1\}^{d}\to\{0,1\}$. For each $i\in[m]$, the $i$-th output bit is obtained by applying $P$ to the input bits on the $i$-th hyperedge. That is, suppose the $i$-th hyperedge contains vertices $v_{i,1},v_{i,2},\dots,v_{i,d}$, then on input $x\in\{0,1\}^{n}$, the $i$-th output bit is